OpenStudy (anonymous):

0How to simplify: ((sqrt54a^3)/a) - (2sqrt24a))

OpenStudy (anonymous):

\[((\sqrt{54a ^{3}})\div a) - (2\sqrt{24a})\]

OpenStudy (anonymous):

sqrt(54a) - 2 sqrt(24a) = 3 sqrt(6a) - 2 sqrt(6a)

OpenStudy (anonymous):

= sqrt(6a)

OpenStudy (anonymous):

does it make sense?

OpenStudy (zzr0ck3r):

its sqrt(54a^3)

OpenStudy (zzr0ck3r):

i c

OpenStudy (zzr0ck3r):

factor out an a^2 and its a on the outside, this cancles the a on bottom

OpenStudy (zzr0ck3r):

sqrt(54a^3) = sqrt(54*a*a^2) = a*sqrt(54a)

OpenStudy (zzr0ck3r):

3 sqrt(6a) - 2 sqrt(6a) is wrong 3 sqrt(6a) - 4 sqrt(6a)

OpenStudy (zzr0ck3r):

2sqrt(24*a) = 2*sqrt(4*6*a) = 2*sqrt(4)*sqrt(6a) = 2*2*sqrt(6a) = 4*sqrt(6a) @akash123